OpenStudy (butterflyhope):

Urgent *will medal*: Can someone explain to me how to express 1 3/5 as a percent? Thanks c:

OpenStudy (anonymous):

Do you mean $$1 \frac{3}{5} $$ or $$ \frac{13}{5} $$

OpenStudy (anonymous):

13/5=(13/5)*100 %=260%

OpenStudy (butterflyhope):

The first one ^-^

OpenStudy (anonymous):